# DART: An Automated End-to-End Object Detection Pipeline with Data Diversification, Open-Vocabulary Bounding Box Annotation, Pseudo-Label Review, and Model Training

### v3

@@ 1 Introduction
In recent years, computer vision has witnessed rapid advancements, revolutionizing various applications through innovative methodologies. A significant development is the emergence of generative AI techniques, exemplified by text-to-image Diffusion models [1, 2, 3, 4, 5]. They establish a robust backbone for generating high-fidelity, contextually rich images, which in turn enables other specialized frameworks [6, 7, 8] including DreamBooth[9]. These frameworks build upon this foundation to enhance the personalization and precision of the generated images, providing more nuanced and versatile controls. Another prominent advancement in generative AI is the large multimodal model (LMM), notably GPT-4 series [10, 11]. LMM leverages the synergy between textual and visual data, significantly enhancing the capabilities of comprehensive semantic understanding and interaction in vision-language tasks. Concurrently, traditional computer vision areas like object detection have also experienced expedited breakthroughs. Open-vocabulary object detection (OVD) models like Grounding DINO [12] have expanded the scope of object recognition, allowing systems to detect and classify objects beyond the constraints of predefined categories. Meanwhile, the YOLO (You Only Look Once) families [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] of real-time object detection systems continue to push the Pareto frontier of speed and accuracy, achieving significant advancement through fast and continuous iterations.
@@
Despite the above-mentioned remarkable progress in image generation, LMM, OVD, and real-time object detection, each technology still has its prominent strengths and weaknesses. Image generation produces high-quality images but can sometimes not be photorealistic. LMMs excel in semantic understanding of images but struggle with directly handling numerical data. OVD can detect a wide range of object categories but is often too slow. In contrast, real-time object detection is fast but limited to predefined categories. If the strengths of these four technologies can be effectively combined and integrated, it will undoubtedly result in a significantly improved outcome. The proposed DART pipeline (Figure 1) aims to achieve this complementary combination of image generation, LMM, OVD, and real-time object detection, leveraging their strengths and compensating for their weaknesses to create a comprehensive and efficient object detection solution from data collection to model evaluation.
@@
Ideally, an object detection application should be capable of detecting any target object both quickly and accurately in diverse environments. However, current technologies do not allow for the simultaneous achievement of open-vocabulary detection, real-time capability, high accuracy, and robustness. The most common approach in the industry is to use real-time object detection to meet the essential requirement for speed while manually collecting and annotating as much data as possible to approach the other three goals [24, 25, 26, 27]. Despite the availability of numerous public datasets, they offer limited assistance in this regard because a majority of specific objects are simply not covered by public datasets. For instance, LVIS [28], one of the most diversified object detection datasets, only includes little more than 1000 entry-level object categories, which is inadequate for many industrial applications. OVD models, on the contrary, are designed to alleviate the object coverage and manual labeling issue. However, they are excessively demanding in terms of computation and memory, making them unsuitable for real-time and on-device applications. On the other hand, the latest image generation tools, such as Stable Diffusion models, are promising alternatives to the inefficient and costly manual collection of more images of the same target object from different scenarios due to their ability for diverse image generation. However, these models lack the ability to mimic the appearance of a specific subject given by visual prompts. Customized frameworks such as DreamBooth could be applied on top for subject-driven image generation instead to address this issue. The generated images can then serve as inputs for OVD to produce bounding boxes, acting as pseudo-labels for training downstream real-time object detection models. However, the results from generative AI can sometimes be overly fanciful, and the quality of bounding boxes produced by OVD can be difficult to guarantee. This necessitates a filtering mechanism to ensure the quality of both generated images and pseudo-labels for the downstream training. This task was previously thought to be exclusive to humans, but now LMMs are well-suited for this task. Although falling short in performing number-related tasks, such as generating bounding box coordinates directly from an input image, they excel in understanding the semantic context of images. As a result, LMM can review the quality of generated images and annotations by solving a designed vision-language task in a c(VQA) manner. To summarize, the foundational advancements in four areas of computer vision collectively form the backbone of the proposed automated end-to-end object detection pipeline DART to fulfill the four objectives of an ideal object detection application. Specifically, DART utilizes subject-driven image generation frameworks for data Diversification to enhance robustness, OVD for ground truth bounding box Annotation to extend object categories without human efforts, LMM for Reviewing generated images and bounding boxes to ensure accuracy and explainability, and Trained real-time object detector as the final output model.
@@
To prove the effectiveness of the DART pipeline for industrial object detection applications and further exploit its potential, we collect a dataset from the internal database of Libherr, a German-Swiss multinational equipment manufacturer. The collected dataset comprises images of Liebherr products covering a wide variety of heavy machinery such as construction machines, earthmoving equipment, deep foundation machines, mining machines, various types of cranes, material handling machines, and more. We apply dataset-specific preprocessing techniques to the raw images to obtain a clean and refined dataset named Liebherr Products (LP).
@@
In this paper, we instantiate the DART using the current best algorithms for core modules and utilize the pipeline on the LP dataset. Specifically, we use DreamBooth [9] with Stable Diffusion XL (SDXL) [5] for data diversification, Grounding DINO [12] for bounding box annotation, GPT-4o [11] for pseudo-label review, InternVL-1.5 [29] for generated image review, and finally YOLOv8 [21] and YOLOv10 [23] families as the final object detector to be deployed. This instantiation results in an exceptional real-time object detector capable of accurately identifying and localizing all object categories with an average precision (AP) of 0.832. Without using DART, the result is a disastrous 0.064. The dramatic difference between object detection with and without DART is vividly illustrated in Figure 13 and Figure 15 through qualitative comparisons. Quantitatively, we also demonstrate the effectiveness of all four major components of DART by showing the continuous improvement in detection performance with the incremental addition of each block. Notably, all four major blocks in the DART pipeline are fully automated, completely eliminating the need for human intervention. This starkly contrasts traditional supervised training methods that require extensive human efforts in labeling and data collection. Additionally, we design the pipeline to be easily interchangeable and extendable, facilitating the adoption of more advanced algorithms in the future and enabling the seamless integration of new data and object categories with minimal human effort.
@@ 2 Related work
In the related work section, we will focus on four areas corresponding to the DART pipeline’s core stages in order: subject-driven image generation, open-vocabulary object detection, and large multimodal models, real-time object detection. By discussing progress in these fields, we aim to justify our current implementation of state-of-the-art methods and explore the potential for further enhancements by incorporating more advanced modules in the future.
@@ 2.1 Subject-driven image generation
Text-to-image diffusion models [1, 2, 3, 4, 5] have swiftly advanced in recent years, establishing themselves as powerful tools in generative AI. These models generate high-quality images from textual descriptions through iterative denoising processes, forming the foundation for subject-driven text-to-image generation. Given only a few (typically 3-5) casually captured images of a specific subject, subject-driven image generation aims to produce customized images with high detail fidelity and variations guided by text prompts. One typical approach is to fine-tune a special prompt token to describe the concept of the specific subject. Textual Inversion [8] shows evidence that a single-word embedding is sufficient for capturing unique concepts. Recent methods introduce additional modules to the subject-driven image generation framework and only update the newly added modules during training while keeping the pre-trained text-to-image frozen. T2I-Adapter [6] aligns pre-trained T2I models with external control signals extracted from trainable CNNs to enhance controllability in text-to-image generation without greatly affecting the original generation ability. IP-Adapter [7] decouples the cross-attention mechanism by separating the cross-attention layers for text features and image features to achieve image prompt capability for the pre-trained text-to-image diffusion models. InstantID [30] follow the path of decoupled cross-attention and leverage a pre-trained face model to detect and extract face ID embedding from the reference facial image to generate ID-preserving images. Although these fine-tuning methods are quick and simple, the generated images often fail to completely resemble the given subject. Therefore, we adopt a full fine-tuning framework, DreamBooth, which directly fine-tunes both the text-to-image backbone with a few images of a specific subject and the token embedding, enabling the generation of diverse high-fidelity images of the specific subjects. To maximize the diversity and fidelity of the generations, we select the powerful SDXL as the backbone for the DreamBooth framework. We also improve the training process to maximize the potential of Dreambooth with SDXL and design diversified prompts to generate new images with diversified contexts.
@@ 2.3 Large multimodal model
Recent advancements in large multimodal models (LMMs) have been characterized by the release of several groundbreaking proprietary systems, including GPT-4v [10], Claude 3 [36], and Gemini [37]. These models have collectively set a new standard for vision-language understanding, demonstrating unprecedented capabilities in combining visual inputs with advanced language processing. Recently, GPT-4o [11] further extends these capabilities with enhanced contextual awareness and multimodal reasoning, pushing the boundaries of LMM by achieving state-of-the-art performance across a wide range of vision-language tasks. Meanwhile, the open-source community has also been actively advancing in this domain. Flamingo [38] utilizes visual and language inputs as prompts, demonstrating exceptional few-shot performance in VQA. LLaVA series [39, 40, 41] and MiniGPT-4 [42] introduce visual instruction tuning to improve the instruction-following ability of LMMs. InternVL family [43, 29] scales up vision foundation models and trains on a curated high-quality dataset to achieve comparable results to GPT-4v among various multimodal benchmarks. Although open-source models have proven adequate for basic semantic understanding tasks such as image review [44, 45], proprietary models, especially GPT-4o, still demonstrate superior performance in handling intricate visual data and sophisticated multimodal interactions, especially when additional visual hints are applied [46]. However, they both struggle with tasks requiring numerical outputs, such as directly generating bounding box coordinates [47]. Therefore, our DART pipeline only leverages GPT-4o/InternVL-1.5 to judge whether a model-generated output is suitable as pseudo-labels/input data for the downstream real-time object detection training while leaving the specialized task of bounding box annotation to OVD. Additionally, to better handle the review task, we craft specialized prompts tailored to the characteristics of the two LMMs, resulting in excellent outcomes.
@@ 2.4 Real-time object detection
Real-time object detection aims to classify and locate objects in real-time with low latency, which is essential numerous industrial applications. This area has witnessed significant advancements in recent years, primarily driven by the iterative development of state-of-the-art YOLO models. Early development of YOLO from v1 to v3 [13, 14, 15] determined the optimal segmentation of the real-time object detection network structure: backbone, neck, and head. YOLOv4 [16] and YOLOv5 [17] changed the backbone to CSPNet [48] and introduced a series of best practices in computer vision regarding training strategy, post-processing, as well as plugin modules. Based on these improvements, YOLOv6 [20] and YOLOv7 [19] are further fine-tuned in label assignment strategy[49] and loss functions [50, 51, 52], and they utilize Repblock [53] and E-ELAN [54] as the core of their backbone respectively. YOLOv8 [21] introduces the C2f building block for effective feature extraction and fusion, along with a comprehensive set of integrated tools, solidifying its position as the go-to choice for real-time object detection. YOLOv9 [22] proposes programmable gradient information (PGI) to refine the training process and a new lightweight network architecture – Generalized Efficient Layer Aggregation Network (GELAN). Finally, the concurrent work YOLOv10 [23] presents an NMS-free training strategy and polishes the mode architecture with holistic efficiency-accuracy-driven design, contributing to its state-of-the-art real-time object detection capabilities. Considering YOLOv8’s ease of use and YOLOv10’s optimal tradeoff between speed and accuracy, this work employs both the YOLOv8 and YOLOv10 families as the real-time object detectors of choice. Although we have selected the best YOLO models, their reliance on predefined object categories restricts their applicability in open scenarios. Therefore, real-time object detection only serves as the last stage in DART, depending on the other three stages to provide high-quality training data.
@@ 3.1 Dataset preprocessing
Class identification. The raw dataset is sourced from Liebherr’s internal digital system, where images of Liebherr’s products are initially roughly categorized in the database. By refining these preliminary classifications with detailed product descriptions, we establish 23 accurately defined classes, as detailed in Section 4.1. We find the precise definition of the class name essential because it highly affects the performance of the following data diversification phase since ambiguous class definition leads to inferior quality of generated data of that class. As a result, each image in the raw dataset is assigned at least one precise class name based on its primary object, and this broad assignment will be used as one of the text prompts for bounding box annotation, as described in Section 3.3.
@@
Deduplication. The initial raw dataset contains many duplicated images. We employ perceptual hashing (pHash) [55] for preliminary deduplication. However, cash tends to misidentify similar images as duplicates, for example, consecutive shots of the same scene from slightly different angles. These near duplicates are manually retained in the dataset but are ensured to appear only in the training set.
@@ 3.2 Data diversification based on DreamBooth with Stable Diffusion XL
Data diversification and augmentation in DART are deeply rooted in the latest image-generation techniques. One of the most popular ones is Stable Diffusion, which represents a series of latent diffusion models that iteratively denoises random Gaussian noise in the latent space of a pre-trained autoencoder using a U-Net architecture [56] conditioned on textual prompts. Stable Diffusion XL (SDXL) introduces several architectural improvements over previous versions of Stable Diffusion, including a larger U-Net backbone, more powerful text encoders, and novel conditioning schemes to enhance the visual fidelity of generated samples, resulting in significantly improved performance. However, SDXL takes only text prompts and, therefore, cannot mimic the appearance of a specific subject given by visual prompts, which is essentially our goal for the data diversification stage. DreamBooth addresses this issue by fine-tuning a pre-trained SDXL model using a few images of a specific subject as ground truth. The key idea is to bind a unique textual identifier with the visual subject, allowing the model to synthesize novel photorealistic images of the subject in diverse scenes and poses with the help of diversified text prompts. The framework of DreamBooth with SDXL is illustrated in Figure 2. Without loss of generality, we showcase the class ”articulated dump truck” and one of its instances, ”TA230”.
@@
Then, the same SDXL model is fine-tuned by setting $\mathbf{\theta}$ and $\mathbf{\theta^{\prime}}$ to be trainable. In the actual implementation, we do not train the original parameters of the SDXL model and its text encoder directly but adopt LoRA (Low-Rank Adaptation) [57] instead which freezes the pre-trained model weights and injects trainable low-rank matrices into the transformer layers’ weight matrices to enable efficient adaptation with minimal additional parameters. Here, we slightly abuse the notation of $\mathbf{\theta}$ and $\mathbf{\theta^{\prime}}$ to also represent LoRA parameters.
@@
where $\alpha_{t}$, and $\sigma_{t}$ are functions of timestep $t$ that control the noise schedule of the diffusion process. As for $w_{t}$ in Equation 1, we adopt Min-SNR weighting strategy [58] to avoid the inefficiency and instability caused by conflicts among timesteps. Specifically, $w_{t}=\min(\alpha_{t}^{2}/\sigma_{t}^{2},\gamma)$, where $\gamma$ is a hyperparameter to avoid the model focusing too much on small noise levels. Note that the encoded instance-specific prompt $\mathbf{c}_{\text{ins}}$ is used for the prediction of the ground truth instance. In contrast to class-specific prompts $\mathbf{P_{\text{cls}}}$ that only provide a general description of a class, instance-specific prompts $\mathbf{P_{\text{ins}}}$ should include both a general class description for better image quality and a unique identifier to build the correspondence to the given visual instance. Therefore, $\mathbf{P_{\text{ins}}}$ utilizes the extracted instance name (usually the product number) wrapped in angle brackets as the identifier (e.g. ”¡TA230¿” in Figure 2) followed by the original class name.
The second component of the overall loss is the class-specific prior preservation loss $L_{pr}$ which is essentially the same SE loss but treats the class images $\mathbf{x_{\text{cls}}}$ as ground truth instead. It is introduced to mitigate the overfitting issue of the standard fine-tuning. With both reconstruction loss $L_{\text{rec}}$ and class-specific prior preservation loss $L_{pr}$, DreamBooth encourages diversity in the generated images while preserving the key visual features of the subject. The total loss $L$ can be expressed as follows:

### v4

@@ 1 Introduction
In recent years, computer vision has witnessed rapid advancements, revolutionizing various applications through innovative methodologies. A significant development is the emergence of generative AI techniques, exemplified by text-to-image Diffusion models [1, 2, 3, 4, 5]. They establish a robust backbone for generating high-fidelity, contextually rich images, which in turn enables other specialized frameworks [6, 7, 8], including DreamBooth[9]. These frameworks build upon this foundation to enhance the personalization and precision of the generated images, providing more nuanced and versatile controls. Another prominent advancement in generative AI is the large multimodal model (LMM), notably GPT-4 series [10, 11]. LMM leverages the synergy between textual and visual data, significantly enhancing the capabilities of comprehensive semantic understanding and interaction in vision-language tasks. Concurrently, traditional computer vision areas like object detection have also experienced expedited breakthroughs. Open-vocabulary object detection (OVD) models like Grounding DINO [12] have expanded the scope of object recognition, allowing systems to detect and classify objects beyond the constraints of predefined categories. Meanwhile, the YOLO (You Only Look Once) families [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] of real-time object detection systems continue to push the Pareto frontier of speed and accuracy, achieving significant advancement through fast and continuous iterations.
@@
Despite the above-mentioned remarkable progress in image generation, LMM, OVD, and real-time object detection, each technology still has its prominent strengths and weaknesses. While the state-of-the-art image generation models produce high-quality outputs, the results may not always exhibit full photorealism. LMMs excel in the semantic understanding of images but struggle with directly handling numerical data. OVD can detect a wide range of object categories, but it is often too slow. In contrast, real-time object detection is fast but limited to predefined categories. If the strengths of these four technologies can be effectively combined and integrated, it will undoubtedly result in a significantly improved outcome. The proposed DART pipeline (Figure 1) aims to achieve this complementary combination of image generation, LMM, OVD, and real-time object detection, leveraging their strengths and compensating for their weaknesses to create a comprehensive and efficient object detection solution from data collection to model evaluation.
@@
Ideally, an object detection application should be capable of detecting any target object both quickly and accurately in diverse environments. However, current technologies do not allow for the simultaneous achievement of open-vocabulary detection, real-time capability, high accuracy, and robustness. The most common approach in the industry is to use real-time object detection to meet the essential requirement for speed while manually collecting and annotating as much data as possible to approach the other three goals [24, 25, 26, 27]. Despite the availability of numerous public datasets, they offer limited assistance in this regard because a majority of specific objects are simply not covered by public datasets. For instance, LVIS [28], one of the most diversified object detection datasets, only includes a little more than 1000 entry-level object categories, which is inadequate for many industrial applications. OVD models, on the contrary, are designed to alleviate the object coverage and manual labeling issue. However, they are excessively demanding in terms of computation and memory, making them unsuitable for real-time and on-device applications. On the other hand, the latest image generation tools, such as Stable Diffusion models, are promising alternatives to the inefficient and costly manual collection of more images of the same target object from different scenarios due to their ability to generate diverse images. However, these models lack the ability to mimic the appearance of a specific subject given by visual prompts. Customized frameworks such as DreamBooth could be applied on top for subject-driven image generation instead to address this issue. The generated images can then serve as inputs for OVD to produce bounding boxes, acting as pseudo-labels for training downstream real-time object detection models. However, the results from generative AI can sometimes be overly fanciful, and the quality of bounding boxes produced by OVD can be difficult to guarantee. This necessitates a filtering mechanism to ensure the quality of both generated images and pseudo-labels for the downstream training. This task was previously thought to be exclusive to humans, but now LMMs are well-suited for this task. Although falling short in performing number-related tasks, such as generating bounding box coordinates directly from an input image, they excel in understanding the semantic context of images. As a result, LMM can review the quality of generated images and annotations by solving a designed vision-language task in a Visual Question Answering(VQA) manner. To summarize, the foundational advancements in four areas of computer vision collectively form the backbone of the proposed automated end-to-end object detection pipeline DART to fulfill the four objectives of an ideal object detection application. Specifically, DART utilizes subject-driven image generation frameworks for data Diversification to enhance robustness, OVD for ground truth bounding box Annotation to extend object categories without human efforts, LMM for Reviewing generated images and bounding boxes to ensure accuracy and explainability, and Trained real-time object detector as the final output model.
@@
To prove the effectiveness of the DART pipeline for industrial object detection applications and further exploit its potential, we collect a dataset from the internal database of Liebherr, a German-Swiss multinational equipment manufacturer. The collected dataset comprises images of Liebherr products covering a wide variety of heavy machinery such as construction machines, earthmoving equipment, deep foundation machines, mining machines, various types of cranes, material handling machines, and more. We apply dataset-specific preprocessing techniques to the raw images to obtain a clean and refined dataset named Liebherr Products (LP).
@@
In this paper, we instantiate the DART using the current best algorithms for core modules and utilize the pipeline on the LP dataset. Specifically, we use DreamBooth [9] with Stable Diffusion XL (SDXL) [5] for data diversification, Grounding DINO [12] for bounding box annotation, GPT-4o [11] for pseudo-label review, InternVL-1.5 [29] for generated image review, and finally YOLOv8 [21] and YOLOv10 [23] families as the final object detector to be deployed. This instantiation results in an exceptional real-time object detector capable of accurately identifying and localizing all object categories with an average precision (AP) of 0.832. Without using DART, the result is a disastrous 0.064. The dramatic difference between object detection with and without DART is vividly illustrated in Figure 13 and Figure 15 through qualitative comparisons. Quantitatively, we also demonstrate the effectiveness of all four major components of DART by showing the continuous improvement in detection performance with the incremental addition of each block. Notably, all four major blocks in the DART pipeline are fully automated, completely eliminating the need for human intervention. This starkly contrasts with traditional supervised training methods that require extensive human efforts in labeling and data collection. Additionally, we design the pipeline to be easily interchangeable and extendable, facilitating the adoption of more advanced algorithms in the future and enabling the seamless integration of new data and object categories with minimal human effort.
@@ 2 Related work
In the related work section, we will focus on four areas corresponding to the DART pipeline’s core stages in order: subject-driven image generation, open-vocabulary object detection, large multimodal models, and real-time object detection. By discussing progress in these fields, we aim to justify our current implementation of state-of-the-art methods and explore the potential for further enhancements by incorporating more advanced modules in the future.
@@ 2.1 Subject-driven image generation
Text-to-image diffusion models [1, 2, 3, 4, 5] have swiftly advanced in recent years, establishing themselves as powerful tools in generative AI. These models generate high-quality images from textual descriptions through iterative denoising processes, forming the foundation for subject-driven text-to-image generation. Given only a few (typically 3-5) casually captured images of a specific subject, subject-driven image generation aims to produce customized images with high detail fidelity and variations guided by text prompts. One typical approach is to fine-tune a special prompt token to describe the concept of the specific subject. Textual Inversion [8] shows evidence that a single-word embedding is sufficient for capturing unique concepts. Recent methods introduce additional modules to the subject-driven image generation framework and only update the newly added modules during training while keeping the pre-trained text-to-image frozen. T2I-Adapter [6] aligns pre-trained T2I models with external control signals extracted from trainable CNNs to enhance controllability in text-to-image generation without greatly affecting the original generation ability. IP-Adapter [7] decouples the cross-attention mechanism by separating the cross-attention layers for text features and image features to achieve image prompt capability for the pre-trained text-to-image diffusion models. InstantID [30] follows the path of decoupled cross-attention and leverages a pre-trained face model to detect and extract face ID embedding from the reference facial image to generate ID-preserving images. Although these fine-tuning methods are quick and simple, the generated images often fail to completely resemble the given subject. Therefore, we adopt a full fine-tuning framework, DreamBooth, which directly fine-tunes both the text-to-image backbone with a few images of a specific subject and the token embedding, enabling the generation of diverse high-fidelity images of the specific subjects. To maximize the diversity and fidelity of the generations, we select the powerful SDXL as the backbone for the DreamBooth framework. We also improve the training process to maximize the potential of Dreambooth with SDXL and design diversified prompts to generate new images with diversified contexts.
@@ 2.3 Large multimodal model
Recent advancements in large multimodal models (LMMs) have been characterized by the release of several groundbreaking proprietary systems, including GPT-4v [10], Claude 3 [36], and Gemini [37]. These models have collectively set a new standard for vision-language understanding, demonstrating unprecedented capabilities in combining visual inputs with advanced language processing. Recently, GPT-4o [11] further extends these capabilities with enhanced contextual awareness and multimodal reasoning, pushing the boundaries of LMM by achieving state-of-the-art performance across a wide range of vision-language tasks. Meanwhile, the open-source community has also been actively advancing in this domain. Flamingo [38] utilizes visual and language inputs as prompts, demonstrating exceptional few-shot performance in VQA. LLaVA series [39, 40, 41] and MiniGPT-4 [42] introduce visual instruction tuning to improve the instruction-following ability of LMMs. InternVL family [43, 29] scales up vision foundation models and trains on a curated high-quality dataset to achieve comparable results to GPT-4v among various multimodal benchmarks. Although open-source models have proven adequate for basic semantic understanding tasks such as image review [44, 45], proprietary models, especially GPT-4o, still demonstrate superior performance in handling intricate visual data and sophisticated multimodal interactions, especially when additional visual hints are applied [46]. However, they both struggle with tasks requiring numerical outputs, such as directly generating bounding box coordinates [47]. Therefore, our DART pipeline only leverages GPT-4o/InternVL-1.5 to judge whether a model-generated output is suitable as pseudo-labels/input data for the downstream real-time object detection training, while leaving the specialized task of bounding box annotation to OVD. Additionally, to better handle the review task, we craft specialized prompts tailored to the characteristics of the two LMMs, resulting in excellent outcomes.
@@ 2.4 Real-time object detection
Real-time object detection aims to classify and locate objects in real-time with low latency, which is essential for numerous industrial applications. This area has witnessed significant advancements in recent years, primarily driven by the iterative development of state-of-the-art YOLO models. Early development of YOLO from v1 to v3 [13, 14, 15] determined the optimal segmentation of the real-time object detection network structure: backbone, neck, and head. YOLOv4 [16] and YOLOv5 [17] changed the backbone to CSPNet [48] and introduced a series of best practices in computer vision regarding training strategy, post-processing, as well as plugin modules. Based on these improvements, YOLOv6 [20] and YOLOv7 [19] are further fine-tuned in label assignment strategy[49] and loss functions [50, 51, 52], and they utilize Repblock [53] and E-ELAN [54] as the core of their backbone respectively. YOLOv8 [21] introduces the C2f building block for effective feature extraction and fusion, along with a comprehensive set of integrated tools, solidifying its position as the go-to choice for real-time object detection. YOLOv9 [22] proposes programmable gradient information (PGI) to refine the training process and a new lightweight network architecture – Generalized Efficient Layer Aggregation Network (GELAN). Finally, the concurrent work YOLOv10 [23] presents an NMS-free training strategy and polishes the model architecture with a holistic efficiency-accuracy-driven design, contributing to its state-of-the-art real-time object detection capabilities. Considering YOLOv8’s ease of use and YOLOv10’s optimal tradeoff between speed and accuracy, this work employs both the YOLOv8 and YOLOv10 families as the real-time object detectors of choice. Although we have selected the best YOLO models, their reliance on predefined object categories restricts their applicability in open scenarios. Therefore, real-time object detection only serves as the last stage in DART, depending on the other three stages to provide high-quality training data.
@@ 3.1 Dataset preprocessing
Class identification. The raw dataset is sourced from Liebherr’s internal digital system, where images of Liebherr’s products are initially roughly categorized in the database. By refining these preliminary classifications with detailed product descriptions, we establish 23 accurately defined classes, as detailed in Section 4.1. We find the precise definition of the class name essential because it highly affects the performance of the following data diversification phase, since an ambiguous class definition leads to inferior quality of the generated data of that class. As a result, each image in the raw dataset is assigned at least one precise class name based on its primary object, and this broad assignment will be used as one of the text prompts for bounding box annotation, as described in Section 3.3.
@@
Deduplication. The initial raw dataset contains many duplicated images. We employ perceptual hashing (pHash) [55] for preliminary deduplication. However, pHash tends to misidentify similar images as duplicates, for example, consecutive shots of the same scene from slightly different angles. These near duplicates are manually retained in the dataset but are ensured to appear only in the training set.
@@ 3.2 Data diversification based on DreamBooth with Stable Diffusion XL
Data diversification and augmentation in DART are deeply rooted in the latest image-generation techniques. One of the most popular ones is Stable Diffusion, which represents a series of latent diffusion models that iteratively denoise random Gaussian noise in the latent space of a pre-trained autoencoder using a U-Net architecture [56] conditioned on textual prompts. Stable Diffusion XL (SDXL) introduces several architectural improvements over previous versions of Stable Diffusion, including a larger U-Net backbone, more powerful text encoders, and novel conditioning schemes to enhance the visual fidelity of generated samples, resulting in significantly improved performance. However, SDXL takes only text prompts and, therefore, cannot mimic the appearance of a specific subject given by visual prompts, which is essentially our goal for the data diversification stage. DreamBooth addresses this issue by fine-tuning a pre-trained SDXL model using a few images of a specific subject as ground truth. The key idea is to bind a unique textual identifier with the visual subject, allowing the model to synthesize novel photorealistic images of the subject in diverse scenes and poses with the help of diversified text prompts. The framework of DreamBooth with SDXL is illustrated in Figure 2. Without loss of generality, we showcase the class ”articulated dump truck” and one of its instances, ”TA230”.
@@
Then, the same SDXL model is fine-tuned by setting $\mathbf{\theta}$ and $\mathbf{\theta^{\prime}}$ to be trainable. In practice, we fine-tune all layers of the SDXL model, including the text encoder, using LoRA (Low-Rank Adaptation) [57]. LoRA freezes the pre-trained weights and inserts trainable low-rank matrices into the transformer layers’ weight matrices, enabling efficient adaptation with minimal additional parameters. Here, we slightly abuse the notation of $\mathbf{\theta}$ and $\mathbf{\theta^{\prime}}$ to also represent LoRA parameters in what follows.
@@
where $\alpha_{t}$, and $\sigma_{t}$ are functions of timestep $t$ that control the noise schedule of the diffusion process. As for $w_{t}$ in Equation 1, we adopt the Min-SNR weighting strategy [58] to avoid the inefficiency and instability caused by conflicts among timesteps. Specifically, $w_{t}=\min(\alpha_{t}^{2}/\sigma_{t}^{2},\gamma)$, where $\gamma$ is a hyperparameter to avoid the model focusing too much on small noise levels. Note that the encoded instance-specific prompt $\mathbf{c}_{\text{ins}}$ is used for the prediction of the ground truth instance. In contrast to class-specific prompts $\mathbf{P_{\text{cls}}}$ that only provide a general description of a class, instance-specific prompts $\mathbf{P_{\text{ins}}}$ should include both a general class description for better image quality and a unique identifier to build the correspondence to the given visual instance. Therefore, $\mathbf{P_{\text{ins}}}$ utilizes the extracted instance name (usually the product number) wrapped in angle brackets as the identifier (e.g., ”¡TA230¿” in Figure 2) followed by the original class name.
The second component of the overall loss is the class-specific prior preservation loss $L_{pr}$, which is essentially the same SE loss but treats the class images $\mathbf{x_{\text{cls}}}$ as ground truth instead. It is introduced to mitigate the overfitting issue of the standard fine-tuning. With both reconstruction loss $L_{\text{rec}}$ and class-specific prior preservation loss $L_{pr}$, DreamBooth encourages diversity in the generated images while preserving the key visual features of the subject. The total loss $L$ can be expressed as follows: